OpenStudy (anonymous):

Multiply and Simplify: √6(√3 + 5√2)

OpenStudy (shivaniits):

so this would go as.. \[\sqrt{18}+5\sqrt{12}\] i just disclosed the brackets here.. which would go on as:- \[3\sqrt{2}+10\sqrt{3}\] that would give you answer as 21.56... hope it works!!

OpenStudy (anonymous):

thanks that did help!! :)

OpenStudy (shivaniits):

glad to hear it helped!!